OpenStudy (anonymous):

What are the x-intercepts of the following quadratics (solve only by factoring): a) x2-10x+6y=-16 b) x2-10x+10y=0 c) x2-10x+12y=11 BEST ANSWER GETS MEDAL!!

OpenStudy (kayders1997):

Do you know how to factor?

OpenStudy (kayders1997):

Are you sure the y's are in these problems?

OpenStudy (anonymous):

This is exactly the problem that was written, yeah.

OpenStudy (kayders1997):

I would know how to factor without the y's this is weird

OpenStudy (anonymous):

Yeah haha

OpenStudy (kayders1997):

Wait an x intercept is when y=0

OpenStudy (kayders1997):

So all the y's would go away in these problems and than you would have to set all of them equal to 0 before you factor

OpenStudy (kayders1997):

Does this make sense?

OpenStudy (anonymous):

I think so

OpenStudy (kayders1997):

Okay so on the first one once you set it equal to 0 you will get?

OpenStudy (kayders1997):

Plug 0 into y because every x intercept is where y=0 so you will have?

OpenStudy (kayders1997):

X^2-10x+6(0)=-16

OpenStudy (fortytherapper):

\[x^2-10x+6y=-16\] That's the first one?

OpenStudy (anonymous):

So i got to the last one, and i'm left with x^2 - 10x = 0

OpenStudy (fortytherapper):

\[x^2-10x+12y=11\] You got: \[x^2-10x = 0\] Is there anyway to factor this? (x )(x )

OpenStudy (anonymous):

not sure

OpenStudy (anonymous):

can one of them be 0?

OpenStudy (fortytherapper):

I don't see a way in factoring it. I would start off like kayders said. Plug in 0 for y, because the x-intercept is where y=0

OpenStudy (anonymous):

i did that,

OpenStudy (anonymous):

it was originally x^2 - 10x + 10y = 0

OpenStudy (fortytherapper):

Oh, this is for b.). My mistake, I thought it was for the last one

OpenStudy (anonymous):

oh my bad sorry i was wrong

OpenStudy (fortytherapper):

Then yes, you can use 0 as one of them. I see it working

OpenStudy (kayders1997):

What happened to the 11?

OpenStudy (kayders1997):

If you have an x in bothparts youcan take out an x(x-10) would be left

OpenStudy (kayders1997):

So x=10

OpenStudy (kayders1997):

And 0

OpenStudy (fortytherapper):

The 11 was for c.) This issue was for b.) And yes, 11 and 0 are the intercepts

OpenStudy (fortytherapper):

I mean 10 and 0